OpenStudy (anonymous):

what would be the polynomial function if the given zeros are -5+i and -5-i?

OpenStudy (anonymous):

is it (-5x+i)(-5x-i)?

mathslover (mathslover):

No agent! If you're given with zeroes : a and b then the polynomial function is : (x-a)(x-b) = 0 Here, you are given that : Zeroes => (-5-i) and (-5 + i ) So, it becomes : (x-(-5-i))(x-(-5+i)) = 0 (x+ 5 + i ) ( x + 5 - i) = 0

mathslover (mathslover):

Multiply : \((x+5+i)(x+5-i) \) and you will get your polynomial function having the zeroes : (-5 + i ) and (-5 - i)